# Demystifying Behavior-Based Malware Detection at Endpoints

### v1

@@ 6.2. Case Studies
Following our measurements, we present two case studies illustrating the sample distribution challenges in endpoint detection.
@@
Difficult Benign Samples. Table 13 shows that Microsoft is the top benign publisher in our EP test set, covering 7.8% of all benign samples. This is counterintuitive as Microsoft is a trusted publisher, and our EP dataset only records samples that could not be classified statically as malware or benign. We discovered that almost all (98%) Microsoft samples in our EP test set share the AM_Delta prefix in their filenames, corresponding to periodic patches to Windows Defender Anti-Malware. The NGR model in EP->EP outputs an average score of 37% on AM_Delta endpoint traces, almost 2.5$\times$ of the average score across all benign traces. Note that the score quantifies the model’s confidence that the input trace belongs to a malware sample. In EP->EP, on a test set containing only AM_Delta samples as the benign samples, the model achieves only 62.6% AUC, compared to 87.5% with all benign samples. Moreover, there are online reports (amdeltareport, ) about anti-malware software falsely flagging AM_Delta files. A popular configuration repository for Sysmon—a tool to monitor process activities—includes an allow-list rule specifically for AM_Delta (amdelta_syzmon, ). Our evidence suggests that AM_Delta samples are difficult to classify correctly due to making sensitive modifications on hosts. Conversely, our SB test set contains only two AM_Delta samples (less than 0.01% of all benign samples). This highlights that an endpoint detector often faces difficult—false positive prone—benign samples in addition to difficult malware families, which is a root cause behind lower performance in SB->EP and EP->EP.
@@
Table 8. Our strategies to improve the performance against sample distribution challenges in endpoint detection.
6.3. Combating Sample Distribution Challenges
Building an intuition from our previous experiments, we propose two strategies to improve the performance in SB->EP and EP->EP.
Soft-Labeling Against Label Noise in EP->EP. Malware detectors are generally trained using hard binary labels—0 for benign and 1 for malware. However, hard labels are hazardous when noisy as they force the model to overfit to a wrong prediction (li2019learning, ). The high frequency of borderline, i.e., potentially noisy, samples exacerbates this problem in endpoint detection. Research suggests that soft labels can be effective against label noise (lukasik2020does, ) by preventing the model from getting too confident on noisy labels during training. Hoping to improve the EP->EP performance, we implement a function to assign soft labels to our borderline training samples.
@@ 6.3. Combating Sample Distribution Challenges
More Accurate Training Distributions in SB->EP. Our models in SB->EP were trained on the distribution of samples captured in public corpora. Although this is standard for obtaining the best SB->SB performance (jindal2019neurlux, ; trizna2023nebula, ), it is not ideal for SB->EP where the test set follows a different distribution. ML models are heavily influenced by the priors encoded in their training sets due to the empirical risk minimization (ERM) paradigm, which minimizes the average loss over all training samples. For example, 10% of the malware in the SB training set is from the Virut family, and, therefore, that family will be prioritized over Emotet, which accounts for 0.3%.
We first attempt to alleviate this discrepancy in SB->EP by resampling the SB training set to include an equal number of samples from each family. Unfortunately, as presented in Table 8 (second segment), this uniform sampling strategy ends up hurting the performance in SB->EP. Considering over 700 malware families are in the SB training set (vs. around 100 in the EP training set), families relevant to endpoint detection remain a minority in the new family-balanced training set. This explains the low performance and raises doubts about whether such an uninformed strategy could work.
Finally, we experiment with an informed strategy where the SB->EP training set is resampled to follow the family distribution in the EP training set. Widely used threat intelligence platforms can guide practitioners about which families to include in training (ciscothreatintel, ). Table 8 (third segment) shows that this strategy brings moderate performance improvements: 3.3% and 1.2% higher TPR for NGR and HYB, respectively. We will continue using this strategy to train models in the rest of our experiments. We leave further improvements to future work, such as handling notably noisy family labels (sebastian2020avclass2, ; kotzias2019mind, ).
7. Environment-Sensitive Program Behaviors Hurt Endpoint Detection
We have shown the performance implications of the distribution of samples encountered by endpoint detectors. However, even when we account for distribution differences (EX4 in Table 7), the performance in SB->SB (63.2% TPR) is higher than the performance in SB->EP (16.7% TPR) and in EP->EP (49.5% TPR). In this section, we aim to illuminate the remaining performance gap between by studying the impact of variable program behaviors.
It is known that program behaviors are environment sensitive and, as a result, a sample can exhibit varying behaviors in different environments (balzarotti2010efficient, ; lindorfer2011detecting, ; avllazagaj2021malware, ). Moreover, malware samples are more environment-sensitive than benign samples because, for example, they execute only when specific vulnerabilities are present in the host (avllazagaj2021malware, ) or do not execute if they detect they are running in a sandbox (lindorfer2011detecting, ; miramirkhani2017spotless, ). In this section, we build on these insights to understand the impact of variable program behaviors on endpoint detection performance. We focus on the variability stemming from environment differences between sandboxes and endpoint hosts (for the SB->EP scenario) and between different endpoint hosts (for the EP->EP scenario). We perform our experiments using the improved NGR models from Section 6.3.
7.1. The Diversity Among Endpoint Traces
Table 9 shows how much traces vary when similar sets of samples (e.g., from the same family) are run in different environments. We aggregate all the traces from three popular malware families and two benign publishers from our SB1, SB2, and EP datasets. To cover different types of samples, we select Wannacry (ransomware), Emotet (banking trojan), and Khalesi (info-stealer) as the families, and Opera (web browser), Roblox (game platform) as the publishers. We then measure the average pairwise distances between these traces within an environment, e.g., between the traces from SB1 (SB1-SB1), and between two environments, e.g., between the traces from SB1 and EP (SB1-EP). We use normalized compression distance metric, which has been used in similar contexts (bailey2007automated, ; avllazagaj2021malware, ).
Within Environment Between Environments SB1-SB1 SB2-SB2 EP-EP SB1-SB2 SB1-EP SB2-EP Wcry 0.16 0.11 0.44 0.28 0.67 0.50 Emot 0.48 0.35 0.50 0.55 0.65 0.53 Khal 0.33 0.29 0.48 0.52 0.62 0.57 Oper 0.05 0.07 0.69 0.43 0.70 0.75 Rblx 0.18 0.04 0.63 0.38 0.67 0.69
Table 9. Average pairwise distances between traces of samples from three malware families and two benign publishers.
We observe: (i) traces from the same sandbox are almost always very similar; (ii) traces from two endpoint hosts are dissimilar (also shown in (avllazagaj2021malware, )); (iii) traces from different sandboxes are dissimilar but much less so than one trace from a sandbox and one trace from an endpoint host. The observation (iii) is potentially detrimental for SB->EP, as the features learned from SB traces might be irrelevant to EP traces, and (ii) is potentially detrimental for both scenarios. The dissimilarity between SB and EP traces implies that the training set of a model in SB->EP follows a different distribution than its test set, which explains the high drift detection in Section 5.1.
Figure 3. Relationship between the average distance of an input EP trace to the SB traces and the model’s output score in SB->EP. Each marker represents the avg. of 30 input traces.
@@ 7.1. The Diversity Among Endpoint Traces
[…]
Figure 4. Prediction score std. deviations on the endpoint traces of each sample in the EP test set; in SB1->EP and EP->EP.
7.2. Case Studies
@@
After our quantitative results on behavior variability and its impact on the model, we present two case studies demonstrating its nature.
@@
Behavior Variability in Benign Samples. RobloxLauncher and OperaPatcher are two benign samples on the opposite ends of the variability spectrum. RobloxLauncher accepts user input through a graphical interface, such as mouse clicks on menu items. We clustered the EP traces of this sample to discover two main execution paths: (i) downloading a long list of assets (such as .mp3 or .jpeg files), presumably to update the game; (ii) creating temporary Internet files and starting a process, presumably for launching the game. Our EP->EP model makes more accurate predictions on the traces in (i) than (ii) (26% vs 38% average prediction score). In its SB2 trace, the sample downloads a similar list of assets, whereas, in its SB1 trace, it creates only a log file and stops, likely because it could not access the Internet. Neither sandbox triggers the execution of the game launch code path, so we expect that a model trained on sandbox traces would struggle to classify endpoint traces that launch the game. Applied to a model trained on SB1 traces, a drift detector (Section 5.1) consistently rejects the endpoint traces of this sample. We believe this illustrates a broader challenge: emulating user interactions in sandboxes is challenging (liu2022enhancing, ), causing major differences in the data distribution of sandbox vs endpoint traces from benign samples. On the other hand, OperaPatcher performs almost the same actions in all its SB1, SB2, and EP traces: creating localization and library files to update the browser. The fact that this sample executes without any user input eliminates behavior (and prediction) variability across traces.
@@
Non-Malicious Malware Traces. In behavior-based detection, most commonly, the label of a sample (e.g., obtained from VirusTotal) is transferred to all traces from this sample, which are then used for training and testing. This practice implicitly assumes that all traces of a malware sample contain some discernible malicious activity. However, this can be problematic in cases where the sample has failed to execute, e.g., because its remote infrastructure is down (yong2021inside, ).
@@
To assess this approach, we study the Wannacry ransomware family. We select Wannacry as it is thoroughly dissected (unlike most families), and its indicators-of-compromise (IOCs) are well-known, allowing us to gauge whether a particular trace is associated with compromise. Using multiple sources, we create a list of IOCs for Wannacry and look for them in our SB1, SB2, and EP traces. We separate the traces into two sets based on whether they contain any IOC and then measure the average prediction scores of our models on each set. Based on Table 14, we observe (i) most traces contain at least one IOC (e.g., 97.7% of the EP traces), (ii) the model’s average prediction score is lower on the traces with no IOC (e.g., 97% vs. 52% in EP->EP); however, it is still significantly higher than the averages on benign traces (13%). The ability to associate malware traces with no IOC with maliciousness is a strength of ML-based methods over IOC-based detectors. Any malware that reaches the execution stage is a potential danger, regardless of whether it successfully compromises the host.
7.3. Sandbox Features That Do Not Generalize
@@
We have shown that sandbox and endpoint traces for a given sample are different. In this section, we examine why they are different and how this difference might hurt the model in the SB->EP scenario.
@@
Sandbox-Specific Features. We found several sandbox-specific features that are highly prevalent (seen in many traces) and predictive (seen only in malware traces), but occur only in one sandbox and not in the other environments. ML models typically exploit such features to minimize the loss and may overlook other predictive features (pezeshki2021gradient, ), causing the model to perform poorly on traces collected from a different sandbox or from endpoints. Table 10 presents four of these features (called artifacts) that we have found in SB1 and SB2 traces. We have identified over 100 such artifacts, most of which frequently co-occur. For each artifact, we report its prevalence (Prv.) and its malware ratio (MalR.) in SB1, SB2, and EP training sets. Prv. is the percentage of samples in which the artifact is present, and MalR. is the percentage of these samples labeled as malware.
@@
File Name SB1 Traces SB2 Traces EP Traces Prv. MalR. Prv. MalR. Prv. MalR. SogouExp. 9.0% 100.0% 0.2% 0.0% PersonalBankP. 3.3% 100.0% Spotify 0.1% 0.0% 3.0% 100% 0.2% 0.0% Python 3.1% 100% 0.1% 0.0%
@@
Table 10. Some strong malware features found in specific sandboxes that do not generalize to other environments.
@@
These artifacts appear to be a result of the specific configuration and particular apps pre-installed on the specific sandbox (see Appendix C). As real-world hosts and other sandboxes do not share the same configuration, this explains why an endpoint classifier trained on traces from one sandbox might fail to generalize to endpoint traces seen in the wild.
The Impact of Sandbox Evasion. Recent works have measured that 40-80% of malware uses at least one evasive technique (galloro2022systematical, ) to avoid analysis or terminate early if it is running in a sandbox (mcafeeevasion, ; miramirkhani2017spotless, ). Evasion makes sandbox traces dissimilar to endpoint traces. Although sandbox evasion is well understood, its implications for endpoint detection have not been measured.
We explore the impact of sandbox evasion. We use a standard heuristic and treat a sample as evasive if the number of actions it performs is too low (kuchler2021does, ; balzarotti2010efficient, ; miramirkhani2017spotless, ). We first find the malware families common between our EP and SB1 test sets (30 total). We then compute the average trace length of each common family using the traces of the samples belonging to it. We count only registry actions, as they can be recorded unambiguously, unlike actions such as process injections, which might have vendor-specific definitions. We then find the length differences between SB1 and EP traces of each family. For example, Wannacry and Gandcrypt have differences of $+11$ and $-6$, respectively. We then split the families into two sets: the 21 families whose SB traces are longer (e.g., Wannacry) are deemed less likely to be evasive, and the 9 families whose SB traces are shorter (e.g., Gandcrypt) are deemed more likely. The median length differences for the non-evasive and evasive families are $+4.4$ and $-4.3$, respectively, with a few outliers on both sides, e.g., $+81$ for Vobfus and $-85$ for Bypassuac. Finally, we measure the TPR of our model in SB1->SB1 and SB1->EP on these two malware sets individually while keeping the benign samples the same.
Table 11. The performance (TPR) on malware families more (Evasive) or less (Non-Eva) likely to be evading sandboxes.
The trace length of a sample monotonically increases as its execution continues. This means our measurements might be confounded by the discrepancies between the execution durations in SB1 and endpoint hosts. Before presenting our results, we confirm this is unlikely to be the case: 82% of our EP traces are from executions that lasted less than 90 seconds. Although we do not know the exact configuration of SB1, allowing one to two minutes of execution is standard for most sandboxes in practice (kuchler2021does, ).
Based on the results in Table 11, we observe: (i) SB1->EP performance is higher (by 3%) on non-evasive families than on evasive families; (ii) SB1->SB1 performance is significantly higher (by 7%) on evasive families than on non-evasive families. This suggests that the classifier exploits short traces (evasiveness) as a feature, which is beneficial for SB->SB, though it does not generalize to SB->EP. However, the non-trivial SB1->EP performance on evasive families might hint that there are still features useful for endpoint detection in the sandbox traces of evasive samples. Although we are limited to observational data, a controlled study on evasive malware to disentangle these features is a promising direction.
These plots support our hypothesis: the model predicts a higher score for the shortest traces in both test sets, i.e., it has a length bias. This bias leads to accurate predictions in the SB set, where very short traces are much more likely to be malware (the malware ratio on the leftmost side of the upper plot is high). However, this bias leads to inaccurate predictions in the EP set, where very short are not more likely to be malware. On the EP traces shorter than 10 actions ($\sim$20of all EP traces), the model achieves 15% TPR (vs. 20% when all traces are kept). Ultimately, trace length is a spurious correlation learned from sandbox traces that fails to generalize to endpoint traces. Evidence suggests this is introduced by evasive malware that tends to produce short sandbox traces in the same execution time. Methods preventing the model from learning such correlations (kirichenko2023last, ; pezeshki2021gradient, ) offer a promising next step.
7.4. Combatting Environment-Sensitivity With Environment-Invariant Features

### v2

@@ VI-B Case Studies
Next, we present two case studies illustrating sample distribution challenges in endpoint malware detection.
@@
Difficult Benign Samples. Table XIV lists Microsoft as the top benign publisher in our EP test set, covering 7.8% of all benign samples. This is counterintuitive as Microsoft is a trusted publisher, and our EP dataset only records samples that could not be classified statically as malware or benign. We discovered that almost all (98%) Microsoft samples in our EP test set share the AM_Delta prefix in their filenames, which correspond to periodic patches to Windows Defender. The NGR model in EP->EP outputs an average score of 37% on AM_Delta endpoint traces, almost 2.5$\times$ of the average score across all benign traces. Note that the score quantifies the model’s confidence that the input trace belongs to a malware sample. In EP->EP, on a test set containing only AM_Delta samples as the benign samples, the model achieves only 62.6% AUC, compared to 87.5% with all benign samples. Moreover, there are reports [65] about anti-malware software falsely flagging AM_Delta files. A popular configuration repository for Sysmon—a tool to monitor process activities—includes an allow-list rule specifically for AM_Delta [66]. These suggest that AM_Delta samples are difficult to classify correctly as they make sensitive modifications on hosts. In contrast, our SB test set contains only two AM_Delta samples (less than 0.01% of all benign samples). This highlights that an endpoint detector often faces difficult—false positive prone—benign samples in addition to difficult malware families, which is a root cause behind poor performance in SB->EP and EP->EP.
@@
VI-C Combating Sample Distribution Challenges
Building on our observations, we propose two strategies to improve the performance in SB->EP and EP->EP.
Soft-Labeling Against Label Noise in EP->EP. Malware detectors are generally trained using hard binary labels—0 for benign and 1 for malware. However, hard labels are hazardous when noisy as they force the model to overfit to a wrong prediction [67]. The high frequency of borderline, i.e., potentially noisy, samples exacerbates this problem in endpoint detection. Research suggests that soft labels can be effective against label noise [29] by preventing the model from getting too confident on noisy labels during training. To improve the EP->EP performance, we implement a function to assign soft labels to our borderline training samples.
[…]
@@ VI-C Combating Sample Distribution Challenges
TABLE VII: Our strategies to improve the performance against sample distribution challenges in endpoint detection.
More Accurate Training Distributions in SB->EP. Our models in SB->EP were trained on the distribution of samples captured in public corpora. Although this is standard for obtaining the best SB->SB performance [6, 7], it is not ideal for SB->EP where the test set follows a different distribution. ML models are heavily influenced by the priors encoded in their training sets because the empirical risk minimization (ERM) paradigm minimizes the average loss over all training samples. For example, 10% of the malware in the SB training set is from the Virut family, which will be prioritized over Emotet, which accounts for 0.3%.
To alleviate this discrepancy in SB->EP, we turn our attention to ML research, where group robustness techniques [31, 69, 70] have been effective against a similar problem. These techniques aim to train models that perform well on each subgroup (e.g., demographic groups) instead of favoring groups over-represented in the training set. We implement a prior group robustness method [71] by creating an SB training set with an equal number of samples from each malware family. Unfortunately, as presented in Table VII (second segment), this uniform sampling strategy ends up hurting the performance in SB->EP. We attribute this shortcoming to the large number of groups (families) in our problem, compared to standardized ML benchmarks, over 600 vs. typically less than 10 [69, 72]. Moreover, malware family labels are often incomplete or noisy [59, 73]. For example, in our SB1 training set, there are many noisy malware family tags such as vbkryjetor (13 samples). As a result, the model is forced to balance hundreds of groups, which also might be inaccurately tagged. These results highlight a key difference between real-world security applications and simplified ML benchmarks, on which most solutions are developed and evaluated, raising doubts about whether an uninformed training strategy could work.
Finally, we experiment with an informed strategy where the SB->EP training set is resampled to follow the family distribution in the EP training set. Widely used threat intelligence platforms can guide practitioners about which families to include in training [74]. Table VII (third segment) shows that this strategy brings moderate improvements: 3.3% and 1.2% higher TPR for NGR and HYB, respectively. We will use this strategy to train models in the rest of our paper.
Takeaways: The distribution of samples selected evaluate malware detectors must account for the detector’s position in a pipeline. Accordingly, creating realistic test sets for ML-based behavioral detectors (i.e., the final step in the pipeline) reveals the overestimated performance in prior evaluations done on broad distributions (63% vs. 93% TPR).
VII Environment-Sensitive Program Behaviors Hurt Endpoint Detection
We have shown the performance implications of the distribution of samples encountered by endpoint detectors. However, even when we accounted for distribution differences (EX4 in Table VI), the performance in SB->SB (63.2% TPR) is higher than the performance in SB->EP (16.7% TPR) and in EP->EP (49.5% TPR). In this section, we aim to illuminate the remaining performance gap by studying the impact of variable program behaviors on ML-based approaches.
It is known that program behaviors are environment sensitive and, as a result, a sample can exhibit varying behaviors in different environments [16, 34, 15]. Malware samples are more environment-sensitive than benign samples because, for example, they execute only if the host has specific software vulnerabilities [15] or do not execute if they detect they are running in a sandbox [34, 21]. Building on these insights, we aim to characterize the impact of variable program behaviors on ML-based detectors. We focus on the variability stemming from environment differences between sandboxes and endpoint hosts (for the SB->EP scenario) and between different endpoint hosts (for the EP->EP scenario). We perform our experiments using the improved NGR models from Section VI-C.
VII-A The Diversity Among Endpoint Traces
Table VIII shows how much traces vary when a set of related samples (e.g., from the same malware family) is run in different environments. We aggregate all the traces from three malware families and two benign publishers in our SB1, SB2, and EP datasets. To cover a range of samples, we select Wannacry (ransomware), Emotet (banking trojan), and Khalesi (info-stealer) as the families, and Opera (web browser), Roblox (game platform) as the publishers. We report the average pairwise distances between these traces within an environment, e.g., between the traces from SB1 (SB1-SB1), and between two environments, e.g., between the traces from SB1 and EP (SB1-EP). We use normalized compression distance, which has been used in similar contexts [75, 15].
Within Envs. Between Envs. SB1-SB1 SB2-SB2 EP-EP SB1-SB2 SB1-EP SB2-EP Wcry 0.16 0.11 0.44 0.28 0.67 0.50 Emot 0.48 0.35 0.50 0.55 0.65 0.53 Khal 0.33 0.29 0.48 0.52 0.62 0.57 Oper 0.05 0.07 0.69 0.43 0.70 0.75 Rblx 0.18 0.04 0.63 0.38 0.67 0.69
TABLE VIII: Avg. pairwise distances between traces of samples from three malware families and two benign publishers.
We observe: (i) traces from the same sandbox are almost always very similar; (ii) traces from two endpoint hosts are dissimilar (also shown in [15]); (iii) traces from different sandboxes are dissimilar but much less so than one trace from a sandbox and one trace from an EP host. The observation (iii) can be detrimental for SB->EP, as the features learned from SB traces might be irrelevant to EP traces, and (ii) is potentially detrimental for both scenarios. The dissimilarity between SB and EP traces implies that a model’s training set in SB->EP follows a different distribution than its test set, explaining the high drift detection in Section V-A.
Figure 2: Relationship between the avg. distance of an input EP trace to the SB traces and the model’s output score in SB->EP. Each marker represents the avg. of 30 input traces.
@@ VII-A The Diversity Among Endpoint Traces
Figure 3: Prediction score std. deviations on the EP traces of each sample in the EP test set; in SB1->EP and EP->EP.
@@
The Impact of Sandbox Configurations. Our experiments use sandbox traces obtained from commercial sandboxes with configurations unknown to us, likely optimized to identify malware from its sandbox behaviors. The configuration differences between sandboxes and endpoint hosts (e.g., operating systems, software versions) lead to the large dissimilarity between SB and EP traces we observed. This raises the question of whether a sandbox can be configured to obtain traces that approximate those from real-world hosts. Training on such sandbox traces will likely result in higher performance in SB->EP. We acknowledge this as a valid possibility, but it remains an open research question in the current landscape. There is no accepted standard to configure sandboxes in such a manner, and some research suggests that it would be challenging to replicate endpoint systems, e.g., simulating user interactions [35] or diverse real-world configurations [22]. Nonetheless, currently, the research community collects training sets mainly from a single sandbox for ML-based detection solutions [55, 5, 76]. We are unable to collect our own sandbox traces with various configurations to analyze how sandbox configurations affect detection performance, as this would cause a disparity between our endpoint and sandbox data (explained in Section IV-A). We highlight this as an important direction for future work.
VII-B Case Studies
@@
Next, we present two case studies to illustrate the nature of behavior variability and its impact on ML models.
@@
Benign Behavior Variability. RobloxLauncher and OperaPatcher are two benign samples on the opposite ends of the behavior variability spectrum. RobloxLauncher accepts user input through a graphical interface, such as mouse clicks on menu items. We clustered the EP traces of this sample to discover two main execution paths: (i) downloading a long list of assets (such as .mp3 or .jpeg files), presumably to update the game; (ii) creating temporary Internet files and starting a process, presumably for launching the game. Our EP->EP model makes more accurate predictions on the traces in (i) than (ii) (26% vs 38% average prediction score). In its SB2 trace, the sample downloads a similar list of assets, whereas, in its SB1 trace, it creates only a log file and stops, likely because it could not access the Internet. Neither sandbox triggers the game launch execution path, so we expect a model trained on sandbox traces to struggle to classify endpoint traces that launch the game. Applied to a model trained on SB1 traces, a drift detector (Section V-A) consistently rejects the EP traces of this sample. We believe this illustrates a broader challenge: emulating user interactions in sandboxes is challenging [35], causing major differences in the data distribution of sandbox vs endpoint traces from benign samples. Conversely, OperaPatcher performs almost the same actions in all its SB1, SB2, and EP traces: creating localization and library files to patch a browser. This sample executing without any user interaction eliminates behavior (and score) variability across traces.
@@
Non-Malicious Malware Traces. In behavioral malware detection, most commonly, the label of a sample (e.g., obtained from VirusTotal) is transferred to all traces from this sample, which are then used for training and testing. This practice implicitly assumes that all traces of a malware sample contain some discernible malicious activity. However, this can be problematic in cases where the sample has refused or failed to execute, e.g., because its remote infrastructure is down [56].
@@
To assess this approach, we study the Wannacry ransomware family. We select Wannacry as it is thoroughly dissected (unlike most families), and its indicators-of-compromise (IOCs) are well-known, allowing us to gauge whether a particular trace is associated with compromise. Using multiple sources, we create a list of IOCs for Wannacry and look for them in our SB1, SB2, and EP traces. We split the traces into two sets based on whether they contain any IOC and then measure the average prediction scores of our models on each set. Based on Table XV, we observe (i) most traces contain at least one IOC (e.g., 97.7% of the EP traces), (ii) the model’s average prediction score is lower on the traces with no IOC (e.g., 97% vs. 52% in EP->EP), which is still significantly higher than the averages on benign traces (13%). The ability to associate malware traces with no IOC with maliciousness is a strength of ML-based methods over IOC-based detectors. Any malware that reaches the execution stage is hazardous, regardless of whether it successfully compromises the host.
VII-C Spurious Features From Sandboxes
@@
We have shown that sandbox and endpoint traces for a given sample are different. Next, we examine why they are different and how this difference may trigger unreliable predictions from the ML models in the SB->EP scenario.
@@
Sandbox-Specific Artifacts. We found several sandbox-specific features that are highly prevalent (seen in many traces) and predictive (seen only in malware traces) but occur only in one sandbox and not in the other environments. ML models typically exploit such features to minimize the loss and may overlook other predictive features [77], causing the model to perform poorly on traces collected from a different sandbox or from endpoint hosts. Table IX presents four of these features (called artifacts) that we have found in SB1 and SB2 traces. We have identified over 100 such artifacts, most frequently co-occurring in traces. For each artifact, we report its prevalence (Prv.) and its malware ratio (MalR.) in SB1, SB2, and EP training sets. Prv. is the percentage of samples in which the artifact is present, and MalR. is the percentage of these samples labeled as malware.
@@
File Name SB1 Traces SB2 Traces EP Traces Prv. MalR. Prv. MalR. Prv. MalR. SogouExp. 9.0% 100.0% 0.0% 0.2% 0.0% PersonalB. 3.3% 100.0% 0.0% 0.0% Spotify 0.1% 0.0% 3.0% 100% 0.2% 0.0% Python 0.0% 3.1% 100% 0.1% 0.0%
@@
TABLE IX: Some strong malware features found in specific sandboxes that do not generalize to other environments.
@@
These artifacts appear to be a result of the specific configuration and particular apps pre-installed on the specific sandbox (see Appendix D). As real-world hosts and other sandboxes do not share the same configuration, this explains why an endpoint classifier trained on traces from one sandbox might fail to generalize to endpoint traces seen in the wild.
The Impact of Sandbox Evasion. Sandbox evasion is a common mechanism malware authors use to make a sample’s sandbox trace dissimilar to its real-world traces. If running in a sandbox, evasive malware attempts to avoid analysis or terminate early [78, 21]. In Appendix E, we present a series of quantitative experiments to understand the implications of sandbox evasion for ML-based detectors. We find that an ML-based detector trained on sandbox traces predicts a higher score of maliciousness for the shortest traces. This length bias boosts the performance in the SB->SB scenario, where very short traces are more likely from malware. The same bias, however, hurts the performance in the SB->EP scenario, where very short traces are not more likely from malware. This makes trace length a spurious correlation learned from sandbox traces that fails to generalize to endpoint traces. Our evidence suggests that this correlation is introduced by evasive malware, which tends to produce short sandbox traces [28].
VII-D Invariant Learning to Counter Environment-Sensitivity